A rectangle is 3 times as long as it is wide. The perimeter is 60 cm. Find the dimensions of the rectangle. Round to the nearest tenth if necessary.
Answer by mananth(16949)

let width be x
length =3x
..
Perimeter = 2L+2W
2x+2*3x= 60
2x+6x=60
8x=60
x=7.5 cm
y= 3x = 3*7.5
=22.5
dimensions are 22.5 by 7.5 cm